OpenStudy (laurendexter):

The function f is defined by f(x)=the sqr root 25-x squared for -5 <(or equal to) x<( or equal to) 5. Find f'(x) Write an equation for the line tangent to the graph of f at x=-3. and 2 other questions

satellite73 (satellite73):

\[f(x)=\sqrt{25-x^2}\]?

OpenStudy (laurendexter):

That is correct

satellite73 (satellite73):

\[\large \frac{d}{dx}[\sqrt{g(x)}]=\frac{g'(x)}{2\sqrt{g(x)}}\] by the chain rule in your case \(g(x)=25-x^2\)

OpenStudy (laurendexter):

so i take the derivative of it and I do that by using the chain rule is what you're saying?

satellite73 (satellite73):

yes

OpenStudy (laurendexter):

okay does this seem right?